A train travels between station 1 and 2, the engineer of the train is instructed to start from rest at station 1 and accelerate uniformly between points A and B, then coast with a uniform velocity between points B and C, and finally accelerate uniformly between points C and D until the train stops at station 2. The distances AB, BC, and CD are all equal, and it takes 5 minutes to travel between the two stations. Assume that the uniform accelerations have the same magnitude, even when they are in opposite directions.
a.) How much of this 5 minute period does the train spend traveling between points A and B?
b.)How much of this 5 minute period does the train spend traveling between points B and C?
c.) How much of this 5 minute period does the train spend traveling between points C and D?
1
Expert's answer
2019-10-21T08:55:49-0400
"AB=BC=CD=d"


AB:

"d=0.5at_1^2\to t_1=\sqrt{\frac{2d}{a}}"

"v=at_1=\sqrt{2ad}"

BC:


"t_2=\frac{d}{v}=\frac{d}{\sqrt{2ad}}=\sqrt{\frac{d}{2a}}"

CD:

"t_3=\sqrt{\frac{2d}{a}}"

Thus,


"T=2t+t+2t=5t=5\ min"

"t=1\ min"

a)


"t_1=2t=2\ min"

b)


"t_2=t=1\ min"

c)


"t_3=2t=2\ min"